OpenStudy (anonymous):

Translate each statement into mathematical symbols! I don't know how to help!! a.Two more than x b. Three times as much as y c.Five less than twice a d. three less than the square of n e. four more than twice as much as x f. Ten more than trice x is fifteen g. Two less than twice the square of a is five less than a. h. Twice as much as y is six more than the cube of y i. The sum of the square of a and b is seven more than twice a and b j. Twice x exceeds y by two more than three times the product of x and y

OpenStudy (anonymous):

Two more than x means 2 exceeds x = x + 2

OpenStudy (anonymous):

b)3 times as much as y means: 3 times of y = 3 multiply y = 3y c) twice a means 2a and 5 less than twice a means: 2a - 5 d) square of n is n^2 and 3 less than n^2 means: n^2 - 3 e) twice of x is 2x and 4 more than means : 2x + 4

OpenStudy (anonymous):

A should be y = x+2, not x = x+2. Thats mathematically impossible even. lol

OpenStudy (anonymous):

f) thrice x is 3x and 10 more than 3x is : 3x + 10 and this is 15 so, 3x + 10 = 15

OpenStudy (anonymous):

This applies to all the answers. b should be x = 3 y, etc

OpenStudy (anonymous):

@soati I have written like this: the sentence goes full to exceeds x then I have only written the value not x = x+3 but = x+3..

OpenStudy (anonymous):

you should be more clear then not to confuse the questioneer!

OpenStudy (anonymous):

I think no one is getting confused but you..

OpenStudy (anonymous):

im just saying, youre not being clear. its like you rush to the answer, you do this a lot. just be more careful.

OpenStudy (anonymous):

Okays @soati ...

OpenStudy (ganpat):

a.Two more than x = x +2 b. Three times as much as y = 3y c.Five less than twice a = 2a - 5 d. three less than the square of n = n2 - 3 e. four more than twice as much as x = 2x +4 f. Ten more than trice x is fifteen = 3x + 10 =15 g. Two less than twice the square of a is five less than a.= 2a2 -2 = a-5 h. Twice as much as y is six more than the cube of y = 2y = y3 + 6 i. The sum of the square of a and b is seven more than twice a and b = = (a + b)2 = 2(a +b) + 7 j. Twice x exceeds y by two more than three times the product of x and y = 2x = y + 3(x * y) +2 This is what i interpret.. let some one else on comment on this...

OpenStudy (anonymous):

@Ganpat ith answer is wrong I think..

OpenStudy (ganpat):

@waterineyes : How do u interpret ??

OpenStudy (anonymous):

Tell me is it: (a + b)^2 or a^2 + b^2???

OpenStudy (ganpat):

oh !! yes.. i guess u r right... it should be... a2 + b2

OpenStudy (anonymous):

Okay... Now it is right...